OpenStudy (vishweshshrimali5):

In a triangle ABC, \(h_1,~h_2,~h_3\) are the lengths of altitudes from A, B and C, which are natural numbers forming an increasing H.P. If \(h_2\) = 24, then possible value(s) of a:b:c is/are ___.

OpenStudy (vishweshshrimali5):



OpenStudy (vishweshshrimali5):



OpenStudy (anonymous):

angles?

OpenStudy (vishweshshrimali5):

Nothing given.

OpenStudy (anonymous):

no no a:b:c

OpenStudy (anonymous):

what do u mean by H.P ?

OpenStudy (vishweshshrimali5):

But, this is given that all the altitudes are natural numbers. @aajugdar, no they are for sides. @BSwan , it means harmonic progression.

OpenStudy (anonymous):

harmonic progression

OpenStudy (anonymous):

ive seen somethinglike this before mmm but that was dealing with vectors let me think if its gonna well with this

OpenStudy (dan815):

HP = inverse of arithmetic series

OpenStudy (vishweshshrimali5):

That is \(\large{\cfrac{1}{h_1} + \cfrac{1}{h_3} = \cfrac{2}{h_2}}\)

OpenStudy (vishweshshrimali5):

Since, \(h_2 = 24\) \(\implies \large{\cfrac{1}{h_1} + \cfrac{1}{h_3} = \cfrac{1}{12}}\)

OpenStudy (anonymous):

use sine law :D

OpenStudy (vishweshshrimali5):

Okay

OpenStudy (anonymous):

you have a diophantine equation now, with \(h_1<h_3\) am i right?

OpenStudy (vishweshshrimali5):

So, by sine law I have, \(\large{\cfrac{\sin{C}}{h_1} = \cfrac{1}{b}}\) \(\implies \large{\cfrac{1}{h_1} = \cfrac{1}{b*h_1}}\) @mukushla, well it is not in the syllabus of the exam I was preparing for. So, we cannot use it.

OpenStudy (anonymous):

you just need to consider restrictions on \(h_1\) and \(h_3\)

ganeshie8 (ganeshie8):

@mukushla can we solve equation of type `ax + by + cxy = d` using diophantine method ?

OpenStudy (vishweshshrimali5):

Sorry guys I have to go. I will post my side of try when I am back. Until then, goo day.

OpenStudy (anonymous):

ok, let us not call it diophantine :-) it's just an equation with natural variables. @vishweshshrimali5 see you later ;) @ganeshie8 oh yeah, do you have the general olution for that type of diophantine?

OpenStudy (dan815):

why do u say 1/h1 + 1/h3 = 2/h2 isnt HP in the form 1/h1 + 1/(h1+d) + 1/(h1+2d)=1/h1 + 1/h2 + 1/ h3

ganeshie8 (ganeshie8):

Wow! thats interesting xD I am familiar with solving equations like ax+by =c using diophantine... but never dealt with an equation involving second order terms... I'll open a new thread :)